# A comprehensive framework for multi-fidelity surrogate modeling with noisy data: a gray-box perspective

### v1

@@ 3.4 Truss model
[…]
@@ 3.4.1 MFSM performance and convergence
Similarly to the previous analytical example, we now investigate the performance of our MFSM and its convergence with respect to the noise-free underlying HF truss model by computing the validation error $\epsilon_{\text{val}}$ for increasing HF experimental design size. More precisely, the HF ED size varies from 5 to 160 data points contaminated with the noise following the prescribed distribution, while the LF ED is fixed in all experiments to 300 data points. The larger amount of LF training data used compared to the previous application is due to the higher dimensionality and complexity of this application. Again, for the computation of $\epsilon_{\text{val}}$, we use $N_{test}=10^{5}$ noise-free HF data points generated using Latin Hypercube Sampling in the input space.
Figure 3: Truss model – Convergence of the validation error $\epsilon_{\text{val}}$ for increasing number of HF training data. Comparison of our MF gray-box with a PCE model trained on HF data only. The dashed lines are the corresponding errors of a PCE model trained on LF data only.
@@ 3.4.2 Confidence and prediction intervals
We now proceed to the construction and evaluation of CIs and PIs for our MF truss model. We set the HF ED size equal to 80 samples, as the previous study demonstrated satisfactory performance at this sample size ($\epsilon_{\text{val}}\approx 1\%$), with only marginal improvement observed when doubling the sample amount. The HF data is contaminated with noise $\varepsilon_{\text{H}}\sim{\mathcal{N}}(0,0.0015)$. Furthermore, the LF ED consists of 300 samples. Similarly to the previous application, $N_{\text{B}}=1,000$ bootstrap replications are performed for the construction of the CIs and PIs. Also, for the computation of the evaluation metrics reported in Table 3, $N_{\text{rep}}=10$ replications are performed, and the test sets for the CI and PI evaluation consist of $N_{\text{t}}=10,000$ data points each, obtained as described in Section 3.3.2.
Table 3: Truss model – Confidence and prediction intervals evaluation, where $\varepsilon_{\text{H}}\sim{\mathcal{N}}(0,0.0015),\\ N_{\text{H}}=80,\,\text{and }N_{\text{L}}=300$
@@ 3.5 Real-world application: aero-servo-elastic simulation of a wind turbine
Each combination of wind speed, turbulence intensity, and shear exponent is used to generate wind time series realizations with $12$ different stochastic seeds for Bladed and $24$ different seeds for FAST. Excluding certain combinations of input parameters that are not realistic results in $4,344$ and $33,480$ simulations for Bladed and FAST respectively. Our HF data set is the mean system response over the $12$ time series from Bladed, thus $362$ data points. Moreover, our LF data set is the mean system response over the $24$ time series from FAST, resulting in a total of $1,395$ data points.
In this case study, both the HF and the LF data are considered noisy, due to the stochasticity in the wind turbine simulations. Our MF framework remains applicable as is.
@@ 3.5.1 MFSM performance and convergence
In this application, we explore the performance of our MFSM for increasing HF experimental design size, equal to $10\%,20\%,...,70\%$ of the total HF data available, while keeping the LF ED fixed to all the available LF data. We compute the validation error $\epsilon_{\text{val}}$ of the MFSM on a test set consisting of the $30\%$ of the HF data which was not used for training any of the MFSMs at each given replication: $N_{\text{test}}=0.3\times 362=109$ data points.
@@
Figure 4: Wind turbine application – Convergence of the validation error $\epsilon_{\text{val}}$ for increasing number of HF training data. Comparison of our MF gray-box with a PCE model trained on HF data only. The dashed lines are the corresponding errors of a PCE model trained on LF data only.
@@ 4 Conclusions
In this paper, we presented a novel and comprehensive framework for multi-fidelity surrogate modeling that effectively handles noisy data and incorporates epistemic uncertainty arising from limited training information. Our regression-based approach aims to emulate the assumed underlying noise-free HF model, and provides accurate predictions and denoising capabilities. It also offers uncertainty estimations with respect to not only the underlying HF model, but also to unseen noise-contaminated HF observations in the form of confidence and prediction intervals respectively, constructed using the bootstrap methodology. The proposed framework is applied in the field of gray-box modeling, where noisy measurements, considered as the high fidelity, are combined with white-box computer simulations, treated as the low fidelity. However, our framework is not limited solely to the grey-box scenario of combining experimental data and computer simulations. Its versatility extends to situations where both the HF and the LF components are experiments or simulations.

### v2

@@ 3.4 Truss model
$A_{1}$ [m2] Lognormal $2\cdot 10^{-3}$ $2\cdot 10^{-4}$
$A_{2}$ [m2] Lognormal $1\cdot 10^{-3}$ $1\cdot 10^{-4}$
@@
For a data set of $1,000$ HF and LF samples, the Pearson correlation coefficient between the HF and LF data is 0.82, while the NRMSE is 1.40.
@@ 3.4.1 MFSM performance and convergence
Similarly to the previous analytical example, we now investigate the performance of our MFSM and its convergence with respect to the noise-free underlying HF truss model by computing the validation error $\epsilon_{\text{val}}$ for increasing HF experimental design size. More precisely, the HF ED size varies from 5 to 160 data points contaminated with the noise following the prescribed distribution, while the LF ED is fixed in all experiments to 300 data points. The larger amount of LF training data used compared to the previous application is due to the higher dimensionality and complexity of this application. Again, for the computation of $\epsilon_{\text{val}}$, we use $N_{test}=10^{5}$ noise-free HF data points generated using Latin Hypercube Sampling in the input space. For the choice of basis in all PCEs involved in this example, we use degree adaptivity from degree 1 up to degree 8.
@@
Figure 12: Truss model – Convergence of the validation error $\epsilon_{\text{val}}$ for increasing number of HF training data. Comparison of our MFSM with a PCE model trained on HF data only. The dashed lines are the corresponding errors of a PCE model trained on LF data only.
@@ 3.4.2 Confidence and prediction intervals
We now proceed to the construction and evaluation of CIs and PIs for our MF truss model. We set the HF ED size equal to 80 samples, as the previous study demonstrated satisfactory performance at this sample size ($\epsilon_{\text{val}}\approx 1\%$), with only marginal improvement observed when doubling the sample amount. The HF data is contaminated with noise $\varepsilon_{\text{H}}\sim{\mathcal{N}}(0,0.0015)$. Furthermore, the LF ED consists of 300 samples. Similarly to the previous application, $N_{\text{B}}=1,000$ bootstrap replications are performed for the construction of the CIs and PIs. Also, for the computation of the evaluation metrics reported in Table 3, $N_{\text{rep}}=10$ replications are performed, and the test sets for the CI and PI evaluation consist of $N_{\text{t}}=10,000$ data points each, obtained as described in the previous example.
Table 3: Truss model – Confidence and prediction intervals evaluation, where $\varepsilon_{\text{H}}\sim{\mathcal{N}}(0,0.0015)$, $N_{\text{H}}=80,\,\text{and }N_{\text{L}}=300$
@@ 3.5 Real-world application: aero-servo-elastic simulation of a wind turbine
Each combination of wind speed, turbulence intensity, and shear exponent was used to generate wind fields. A wind field was used as input to the wind turbine simulators to produce a time series realization of the structural response. Due to the stochasticity in the wind field generation, $12$ different stochastic seeds were used for Bladed and $24$ different seeds for FAST. Excluding certain combinations of input parameters that are not realistic resulted in $4,344$ and $33,480$ simulations for Bladed and FAST respectively. Our HF data set is the mean of the output maximum flapwise bending moment at the wind turbine blade root over the $12$ time series from Bladed, thus $362$ data points. Moreover, our LF data set is the mean system response over the $24$ time series from FAST, resulting in a total of $1,395$ data points. There are 261 common input samples between the HF and LF data sets. For these samples, the Pearson correlation coefficient between the HF and LF data is 0.95, with a NRMSE of 0.38.
Both FAST and Bladed are deterministic simulators, meaning that repeated runs with the same initial conditions and same input wind field produce the same output time series. However, since the wind fields are stochastic with respect to the three input random variables, both the HF and the LF data coming from the wind turbine simulators contain stochasticity. We treat this stochasticity as additive homoscedastic noise, and our MF framework remains applicable as is.
@@ 3.5.1 MFSM performance and convergence
In this application, we explore the performance of our MFSM for increasing HF experimental design size, equal to $10\%,20\%,...,70\%$ of the total HF data available, while keeping the LF ED fixed to all the available LF data. We compute the validation error $\epsilon_{\text{val}}$ of the MFSM on a test set consisting of the $30\%$ of the HF data which was not used for training any of the MFSMs at each given replication: $N_{\text{test}}=0.3\times 362=109$ data points. For the choice of basis in all PCEs involved in this application, we use degree adaptivity from degree 1 up to degree 10.
@@
Figure 14: Wind turbine application – Convergence of the validation error $\epsilon_{\text{val}}$ for increasing number of HF training data. Comparison of our MFSM with a PCE model trained on HF data only. The dashed lines are the corresponding errors of a PCE model trained on LF data only.
@@ 4 Conclusions
In this paper, we presented a novel and comprehensive framework for multi-fidelity surrogate modeling that effectively handles noisy data and incorporates epistemic uncertainty arising from limited training information. Our regression-based approach aims to emulate the assumed underlying noise-free HF model, and provides accurate predictions and denoising capabilities. It also offers uncertainty estimations with respect to not only the underlying HF model, but also to unseen noise-contaminated HF observations in the form of confidence and prediction intervals respectively, constructed using the bootstrap methodology. The proposed framework is applied to combine experimental data and computer simulations, where noisy measurements, considered as the high fidelity, are combined with white-box computer simulations, treated as the low fidelity. However, our framework is not limited solely to this particular scenario. Its versatility extends to situations where both the HF and the LF components are experiments or simulations.